OpenStudy (anonymous):

(I'll fan and give you a medal! ) 4. Suppose soda is on sale for $0.50 a can, and have a coupon for $0.80 off your total purchase. Write a function rule for the cost of n cans of soda. A. C(n)=0.80n-0.50 B. C(n)= 0.50n-0.80 c. C(n)= (0.50+n)-0.80 d. C(n)=(0.80+n) -0.50

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@Colere252346

OpenStudy (anonymous):

Do you have any ideas about this?

OpenStudy (bohotness):

here

OpenStudy (bohotness):

you're going to add love

OpenStudy (anonymous):

someone please I'm desperate

OpenStudy (bohotness):

yea

OpenStudy (anonymous):

@Mimi_Styles You have $0.80 off of your purchase, so what will you be subtracting?

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

*cries*

OpenStudy (anonymous):

@AnswerMyQuestions < < < < < !! ! !! ! ! D:

OpenStudy (bohotness):

i am having a liitle lag

OpenStudy (anonymous):

If you think about it, total cost would be $0.50 for every can then minus $0.80 because of the coupon. Let every can be 'x'.

OpenStudy (bohotness):

i think you asnwer might me b or c

OpenStudy (anonymous):

...

OpenStudy (anonymous):

wait huh?

OpenStudy (anonymous):

I should have said let every can be 'n', instead. The function is C(n) = 0.50*n - 0.80

OpenStudy (anonymous):

so it would be C? :o

OpenStudy (anonymous):

where did you go? please come back! I really need to get this done so I can work on my other work :(

OpenStudy (anonymous):

today is the last day! @aleroth

OpenStudy (anonymous):

To try out which one makes sense you could try plugging in numbers for n. Let's try buying 2 cans. The price would be: 0.5 dollars/can * 2 cans - 0.80 dollar coupon = 1 - 0.8 = 0.2 dollars Let's try answer C: C(n)= (0.50+n)-0.80 C(2) = (0.50+2)-0.80 = 2.5-0.8 = 1.7 dollars <-- this is not the same as above, you see? Let's try answer B: C(n)= (0.50*n)-0.80 C(2)= (0.50*2)-0.80 = 1-0.8 = 0.2 dollars <-- This is it!

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

please @AnswerMyQuestions I really need to get this done :(

OpenStudy (anonymous):

A function is a relation in where every input has only one output. So if there are the same outputs, then it is not a function.

OpenStudy (anonymous):

(The outputs here are what the numbers are pointing to)

OpenStudy (anonymous):

how do I figure out which ones are the same outputs?

OpenStudy (anonymous):

Numbers can't be pointing to the same number, or it is not a function, so what is your answer?

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

yay! can you please please pretty please help me with more? I'm stressing over here and I only have until 11:59pm to get this all done!

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

It's 12:12 here :P

OpenStudy (anonymous):

AH!

OpenStudy (anonymous):

it's 12:13pm here

OpenStudy (anonymous):

do you see my next question D:

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

I don't get it..

OpenStudy (anonymous):

"An arithmetic sequence is a sequence with the difference between two consecutive terms constant. The difference is called the common difference."

OpenStudy (anonymous):

what about that one ^ @AnswerMyQuestions

OpenStudy (anonymous):

@Directrix Can you help? I'm a little busy right now.

OpenStudy (anonymous):

*stresses*

OpenStudy (anonymous):

Exercise 6 is an arithmetic sequence with the common difference 6. Which means subtracting two following numbers will result in 6.

OpenStudy (anonymous):

*cries*

OpenStudy (anonymous):

I knew it!

OpenStudy (anonymous):

I really need help on 8 tho D:

OpenStudy (anonymous):

@Directrix HELLOOO D: come back :(

OpenStudy (anonymous):

someone come back :(

OpenStudy (anonymous):

@Mimi_Styles Ask 1 question per post.

OpenStudy (anonymous):

oki.

OpenStudy (anonymous):

Open a new post to ask questions.